### v2

# Photonuclear Cross Sections for the 197Au(,pn)195mPt Reaction Near Threshold

Abstract
Platinum radioisotopes are of growing interest for targeted cancer therapy and diagnostic imaging because their decay delivers highly localized radiation doses in tissue, herewith enabling precise DNA damage through Auger-electron emission. Developing production technologies that provide platinum isotopes with high specific activity is therefore essential, and photonuclear reactions on stable nuclei offer a viable accelerator-based route when supported by reliable cross-section data. We report photonuclear cross-section measurements for the 197Au(,pn)195mPt reaction at incident -ray energies of 27, 29, and 31 MeV using the activation method. The measurements were performed by irradiating a stack of concentric-ring gold targets with a quasi-monoenergetic -ray beam provided by the High Intensity Gamma-ray Source (HIS). The induced 195mPt activity was quantified using off-line -ray spectroscopy. These data provide the first experimental constraints on the 197Au(,pn)195mPt cross section in the near-threshold region. The measured excitation function is compared with PHITS and TALYS calculations and indicates that the reaction becomes measurable only near 30 MeV, with substantially higher bremsstrahlung end-point energies required for practically meaningful production.
I Introduction
Platinum-195m (195mPt) is a metastable radionuclide with a half-life of about 4 days that decays via isomeric transitions. Because of its highly internally converted decay, 195mPt emits a large number of low-energy Auger electrons, which deposit their energy over nanometer-scale distances and have attracted interest for targeted Auger-electron therapy and theranostic applicationsΒ [6, 9]. In addition, the emission of a 98.9-keV -ray makes 195mPt suitable for quantitative SPECT imagingΒ [1]. The medical relevance of 195mPt is closely linked to platinum-based chemotherapy. Platinum compounds, such as cisplatin, are widely used anticancer drugs, and radiolabeling these agents with 195mPt enables direct imaging of their biodistribution and pharmacokinetics without altering their chemical structure. Preclinical studies have demonstrated the feasibility of SPECT imaging and quantitative analysis using 195mPt-labeled compoundsΒ [1, 3], and more recent translational work has further highlighted the potential of 195mPt for imaging chemotherapy delivery in vivoΒ [5].
Production of 195mPt has traditionally relied on reactor-based routes involving neutron capture on enriched iridium targetsΒ [10]. Although these methods can provide substantial yields, they may be limited in achievable specific activity due to the presence of stable platinum isotopes. Accelerator-based production using charged particles, such as the 192Os(,n)195mPt reaction, has been shown to improve specific activity, but typically requires enriched targets and is constrained by available beam currentsΒ [4]. Production of 195mPt via photonuclear reactions using electron accelerators has received comparatively little experimental attention [2]. This approach is attractive because photonuclear reactions can produce the desired radionuclide in a chemically distinct target material, herewith enabling efficient chemical separation and potentially high specific activity, an important requirement for Auger-electron-emitting radionuclides [9]. For feasibility studies of photonuclear production of 195mPt, reliable cross-section data for the relevant reactions are required. For 195mPt-based radiopharmaceuticals, the effective specific activity is governed, not only by the total platinum content, but also by the isomer-to-ground-state production ratio. Theoretical predictions based on TALYS [8] indicate that this ratio remains 1 for photon energies above 30 MeV.
The present work focuses on the production of 195mPt via the 197Au(,pn)195mPt reaction. The metastable isomer emits a characteristic 98.9-keV -ray with a half-life of 4.01 days, allowing selective identification of the reaction channel using off-line -ray spectroscopy. Photonuclear reaction rates were determined using an activation method, as described in Sec. II and III. The measured cross sections are presented in Sec. IV and the feasibility of producing 195mPt via photonuclear reactions is discussed in Sec. V.
II Method
When a photon beam irradiates 197Au, radioactive nuclides are produced via photonuclear reactions, including 197Au(,pn)195mPt and 197Au(,2n)195Au, which both decay with a branch emitting the 98.9-keV -ray. The observed counts of this 98.9-keV peak contain contributions from the decays of both 195mPt and 195Au. The simplified decay schemes for 195mPt and 195Au are presented in Fig. 1. The relationship between the observed -ray counts and the number of radioactive nuclei can be written as
where denotes the observed -ray counts and is the number of radioactive nuclei at the end of bombardment (EOB). Here, is the overall correction factor, which accounts for detector efficiency, detector dead time, branching ratio, -ray attenuation in the target, and self-absorption effects. The symbols , , and denote the decay constant, cooling time, and measurement time, respectively. Because the observed -ray counts represent a superposition of contributions from the two nuclides, measurements performed at different cooling times allow the individual contributions from 195Au and 195mPt to be separated.
In this work, three measurements were performed using different cooling times: a few hours after the end of the bombardment (EOB), 1 week, and 2 weeks later. The corrected -ray counts for -th measurement can be expressed as
where labels the nuclides (=1 for 195Au and =2 for 195mPt); represents the corrected number of decay events corresponding to the observed -ray counts after applying the overall correction factor . The and symbols are the cooling time and measurement time for the -th measurement, respectively. For convenience, the above expression can be rewritten in linear form as
with
The parameters and were determined by a weighted least-squares fit (see Appendix A for details). The solution can be written explicitly as
Statistical uncertainties of and are given by
III Experiment
A stack of five targets, each divided into three segments [12], was irradiated at seven different central photon beam energies at the High Intensity Gamma-ray Source (HIS) . Owing to the radial energy gradient across the target stack, the photon energy at the outermost ring was typically close to the central energy of the subsequent lower-energy irradiation, providing an internal consistency check for the measurements (see ref. [12] for details). Using targets composed of three concentric segments, cross sections for the 197Au(,pn)195mPt reaction were obtained at a total of 3 distinct photon energies for the 31 MeV run, as the cross sections are extremely low below 30 MeV. Each irradiation was performed for 8β12 hours using a photon beam with fluxes in the range of - . Detailed information on incident beam flux, irradiation time, target thickness, and beam energy and flux at each segment is summarized in Tables III, IV, and VIII of Ref. [12]. The measurement and cooling times for 31 MeV run are summarized in Table 1.
| seg. 1 | seg. 2 | seg. 3 | ||||
| (MeV) | 31.05 | 29.25 | 27.24 | |||
| run | hr | hr | hr | hr | hr | hr |
| 1 | 5 | - | 5 | - | 10 | - |
| 2 | 7 | 146.29 | 13 | 151.58 | 12 | 188.19 |
| 3 | 24 | 310.59 | 18 | 332.03 | 24 | 353.69 |
The -ray spectra measured for segment 1, the innermost segment, at three different times are presented in Fig. 2 following irradiation: a few hours after the end of bombardment (EOB), one week, and two weeks later. The data accumulation times for the three measurements were 5, 7, and 24 h, respectively. The full spectrum measured at the earliest time (see Fig. 2 (a)) exhibits multiple -ray lines originating from activation products such as 196g,196m1,196m2Au and 194Au in the gold target. Among these, the -ray line at 98.9 keV is of particular interest, as it is emitted by both 195mPt and 195Au, whose individual contributions cannot be distinguished using only this single measurement.
The expanded spectra around the 98.9-keV region (see Figs. 2 (b)β(d)) demonstrate a clear evolution of the relative peak intensities with cooling time. At early times after irradiation, the spectrum is dominated by contributions from shorter-lived nuclides, whereas measurements performed after longer cooling times increasingly emphasize the decay of longer-lived reaction products. This temporal variation arises from the different half-lives of 195mPt ( d) and 195Au ( d), and enables their individual activities to be disentangled through measurements performed at multiple cooling times. The systematic change in the 98.9-keV peak intensity as a function of cooling time provides the experimental basis of the approach for the activity decomposition described in Sec. II. By combining the spectra measured at different cooling times with a least-squares analysis, the initial activities of the contributing nuclides can be extracted reliably, despite the spectral overlap at this 98.9-keV -ray energy.
IV Results
The photonuclear cross sections corresponding to the number of radioactive nuclei at EOB, and can be expressed as
Here , and represent the number of target nuclei, incident photon flux, and irradiation time, respectively. The statistical uncertainty is evaluated using Eq. 5. Three sources of systematic uncertainty arising from the -ray branching ratio, detector efficiency and beam flux are considered. The total systematic uncertainty can therefore be expressed as
where , , and denote the uncertainties associated with photon beam flux, HPGe detector efficiency, and the branching ratio, respectively. The uncertainties in the photon beam flux and HPGe detector efficiency are 5 and 8, while uncertainties of the branching ratios are given in Table 2.
| Nucl. | decay mode | half-life | energy | branching ratio |
| d | keV | % | ||
| 195mPt | IT 100 | 4.010 (5) | 98.9 | 11.7 (8) |
| 195Au | 100 | 186.01 (6) | 98.9 | 11.21 (15) |
The excitation function of the 197Au(,pn)195mPt measured at incident -ray energies of 27, 29, and 31 MeV is illustrated in Fig. 3. The measured cross sections exhibit a clear increase with photon energy, reflecting the opening and growing contribution of the (,pn) reaction channel above threshold. Vertical error bars represent the total experimental uncertainties, which is the combined statistical (from Eq. 5) and systematic (from Eq. 7) uncertainties. The statistical uncertainty of the observed -ray counts is the dominant contribution to the overall uncertainty in extracting cross sections. This is primarily due to the relatively small contribution of the 195mPt decay to the mixed 98.9-keV -ray peak compared to that of 195Au, with the 195mPt activity accounting for only about 5-18 of the total activity at EOB. For each segment, the measured 195mPt activity at EOB ranges from 4 to 1.2 Ci. As a result, the counting statistics for 195mPt are limited and strong anti-correlation arises between the fitted activities. Consquently, the relative uncertainty of the extracted 195mPt activity is comparable to, or larger than, that of the measured cross section.
The measured cross sections are compared with theoretical predictions calculated using the PHITS [11] and TALYS [8] codes. Our data are consistent with the overall trend of both models within the experimental uncertainties. These results provide the first experimental constraints on the cross section for the 197Au(,pn)195mPt reaction in the energy region just above reaction threshold.
V Conclusion
We have measured the photonuclear cross sections for the 197Au(,pn)195mPt reaction at incident energies of 27, 29, and 31 MeV using a quasi-monoenergetic -ray beam at HIS. The observed 98.9-keV -ray peak contains contributions from the decays of both 195Au and 195mPt. These contributions were successfully extracted using a least-squares analysis of measurements performed at multiple cooling times. For the first time, cross sections were experimentally determined in the vicinity of the reaction threshold. No statistically significant 98.9-keV peak attributable to 195mPt was observed for irradiation energies below 27 MeV, despite the reaction threshold energy of Eth = 13.96 MeV, indicating that, under present experimental conditions, specifically a photon flux of and irradiation time 8-10 hours, the production yield is insufficient to detect this rare reaction channel at lower energies. The measured cross sections indicate that the 197Au(,pn)195mPt reaction begins to contribute measurably at incident photon energies of 30 MeV, and provide the first experimental constraints on the cross section in the near-threshold region. The measured cross sections are consistent, within the experimental uncertainties, with the overall trends predicted by the PHITS and TALYS calculations. Although electron energies 30β40 MeV are commonly employed for photonuclear isotope production, the present results suggest that such energies are insufficient to yield meaningful 195mPt production under realistic experimental conditions. Combining our results with the model predictions for the excitation function for the 197Au(,pn)195mPt reaction shown in Fig. 3, we project that efficient production of 195mPt via photonuclear reactions in quantities required for applications will require bremsstrahlung photon beams with substantially higher end-point energies than are standard, e.g., on the order of 50-60 MeV. The present results provide new benchmark data for calculations of photonuclear production of 195mPt and motivate future measurements at higher -ray energies, extending beyond 50 MeV at HIS.
Acknowledgments
This research was supported by the U.S. Department of Energy Isotope Program, managed by the Office of Science for Isotope RD and Production, under contract DE-AC02-06CH11357 (Argonne National Laboratory), and the Department of Energy, Office of Nuclear Physics under grant numbers DE-SC0018112, DE-SC0018325, DE-FG02-97ER41033 (TUNL) and DEFG02-97ER41041 (UNC). The authors would like to thank the HIS operators and staff for their help during the experiments. We gratefully acknowledge the computing resources provided on Bebop, a high-performance computing cluster operated by the Laboratory Computing Resource Center at Argonne National Laboratory.
Appendix A Weighted least-squares formalism
From Eq. 2, the measured counts can be written as
with
The parameters and were determined by a weighted least-squares fit. The is defined as
where denotes the statistical uncertainty of . The is minimized by solving,
Minimization of this expression leads a set of normal equations that can be written in matrix form as
The solution can be written explicitly as
where .
The statistical uncertainty is calculated as,